OpenStudy (anonymous):

i will medal Which ratio correctly compares 15 m to 45 cm? A. 10 : 3 B. 1 : 3 C. 100 : 3 D. 1 : 300

OpenStudy (solomonzelman):

which part of the problem troubles you?

OpenStudy (anonymous):

all of it @SolomonZelman

OpenStudy (solomonzelman):

okay, do you mind if I show you a similar example?

OpenStudy (anonymous):

i really just need the anwer

OpenStudy (solomonzelman):

We don't give answers like this on this website. 1) The openstudy code of conduct: http://openstudy.com/code-of-conduct 2) Once you get how to do these problems, you won't need anyone's help.

OpenStudy (anonymous):

oh sorry i havent read that befo so could u help me tho

OpenStudy (solomonzelman):

Yes, sure, so I will start with an example:

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

` Which ratio correctly compares 10 m to 50 cm? ` First you would need to convert 10 meters to centimeters. (Provided that 1m=100cm) \(\large\color{black}{ 10~{\rm m} \times\frac{\LARGE 100~{\rm cm}}{\LARGE 1 {\rm m}} ~~~ \Rightarrow~~~ 10~{\rm \cancel{m}} \times\frac{\LARGE 100~{\rm cm}}{\LARGE 1 {\rm \cancel{m}}} ~~~ \Rightarrow~~~ 1000~{\rm cm} }\) So now we can re-write the question as: ` Which ratio correctly compares 1000cm to 50 cm? ` \(\large\color{black}{ 1000~{\rm cm}:50~{\rm cm} }\) \(\large\color{black}{ 1000:50 }\) \(\large\color{black}{ 100:5 }\) ANSWER: \(\large\color{black}{ 20:1 }\)

OpenStudy (solomonzelman):

it took long for me to put an example, sorry.

OpenStudy (anonymous):

I STILL DONT GET IT

OpenStudy (solomonzelman):

you read my example in 2 seconds? very good reading skills, but please take another shut. I will be back in a minute.

OpenStudy (dtan5457):

@makkaerya This is not hard if you listened. Instead of spamming other questions with your own, learn this concept with effort, pleae.

OpenStudy (dtan5457):

*please

OpenStudy (solomonzelman):

you basically need to convert all the units of length to the same unit. I would suggest to convert 15 meters to centimeters, and not 45cm to meters (even though meter is the SI unit)